On Tuesday, the price and quantity demanded of bananas were 10 taka per banana and 100 bananas, respectively. The previous Saturday, the price and quantity demanded were 6 taka per banana and 150 bananas, respectively.  

a) Use the midpoint method to find the PED. 

b) What will happen to the producer’s revenue if the price now rises? Why? 

c) Which effect is stronger on the revenue? Price effect, quantity effect or neither?


1
Expert's answer
2021-04-30T07:24:34-0400

a)

"Q1=150"

"Q2=100"

"P1=6"

"P2=10"


"PED= \frac{percentage \ change\ in\ quantity}{percentage\ change\ in\ price}"


"Percentage \ change\ in \ quantity= \frac{Q2-Q1}{(Q2+Q1)/2} \times100"


"=\frac{100-150}{(100+150)/2} \times 100"

"=-40"

"Percentage \ change\ in\ price= \frac{P2-P1}{(P2+P1)/2}\times100"


"=\frac{10-6}{(10+6)/2}\times100"


"=50"

"PED= \frac{-40}{50}=\frac{-4}{5}" "or\ -0.8"

The elasticity of demand is smaller than one which means that demand in this interval is inelastic.


b) When the price of an inelastic good is increased while demand remains constant, the overall revenue rises as a result of the higher price and constant quantity demanded. This means that businesses selling inelastic products or services will raise prices, resulting in fewer sales but higher profits.


c) A quantity effect occurs when a price is raised and fewer units are sold, resulting in lower revenue. If a good's demand is elastic (i.e., the price elasticity of demand is greater than 1), a price increase decreases gross revenue. The quantity effect is stronger than the price effect in this situation.